Suppose your firm has decided to use a divisional WACO approach to analyze projects. The firm currently has four divisions. A through D with weighted average betas for each division of 0.8,1.2,1.4,and 1.6 respectively. Assume all current and future projects will be financed with 30 percent debt and 70 percent equity,the current cost of equity(based on the firms beta of 1.1 and a current risk free rate of 6 percent) is 13 percent and the after tax yield on the company's bonds is 11 percent. What will the WACO be for each division?
Answer #1

cost of equity = risk free rate + beta * market risk premium

=>

market risk premium = (0.13-0.06)/1.1

= 6.36%

For division A

cost of equity = 0.06 + 0.8 * 6.36%

= 11.088%

WACO = 0.3 * 11% + 0.7 * 0.11088

= 11.06%

For division B

cost of equity = 0.06 + 1.2 * 6.36%

= 13.632%

WACO = 0.3 * 0.11 + 0.7 * 0.13632

= 12.84%

For division C

cost of equity = 0.06 + 1.4 * 6.36%

= 14.904%

WACO = 0.3 * 0.11 + 0.7 * 0.14904

= 13.73%

For division D

cost of equity = 0.06 + 1.6 * 6.36%

= 16.176%

WACO = 0.3 * 0.11 + 0.7 * 0.16176

= 14.62%